ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 164, 2008

Bodin, A. Reducibility of rational functions in several variables  333

Cadoret, A. Lifting results for rational points on Hurwitz moduli spaces

19

Carocca, A., V. González-Aguilera, R. A. Hidalgo and R. E. Rodríguez Generalized Humbert curves

165

Diem, C. and G. Frey Non-constant curves of genus 2 with infinite pro-Galois covers

193

Elin, M., S. Reich and D. Shoikhet A Julia–Carathéodory theorem for hyperbolically monotone mappings in the Hilbert ball

397

Frey, G. See C. Diem

Frohmader, A. Face vectors of flag complexes  153

Glasner, E., M. Megrelishvili and V. V. Uspenskij On metrizable enveloping semigroups

317

González-Aguilera, V. See A. Carocca

Grushevsky, S. and R. Salvati Manni Jacobians with a vanishing theta-null in genus 4

303

Gutman, Y. and M. Hochman On processes which cannot be distinguished by finite observation

265

Han, P. Many solutions for elliptic equations with critical exponents  125

Hidalgo, R. A. See A. Carocca

Hochman, M. See Y. Gutman

Katz, M. Systolic inequalities and Massey products in simply-con-nected manifolds

381

Klartag, B. and E. Milman On volume distribution in 2-convex bodies

221

König, B., P. Larson, J. T. Moore and B. Veličković Bounding the consistency strength of a five element linear basis

1

Larson, P. See B. König

Li, X., Y. Li and L. Wang ℨ-permutable subgroups and p-nilpotency of finite groups II

75

Li, Y. See X. Li

Magyar, A. On distance sets of large sets of integer points  251

Máthé, A. Hausdorff measures of different dimensions are not Borel isomorphic

285

Megrelishvili, M. See E. Glasner Mendelson, S. and N. Tomczak-Jaegermann A subgaussian em-

bedding theorem

349

Milman, E. See B. Klartag Moore, J. T. See B. König Muić, G. On the structure of theta lifts of discrete series for dual pairs

(Sp(n), O(V)) 87

Natapov, M. Projective bases of division algebras and groups of central type II

61

Reich, S. See M. Elin Rodríguez, R. E. See A. Carocca Salvati Manni, R. See S. Grushevsky Shoikhet, D. See M. Elin

Sivatski, A. S. On indecomposable algebras of exponent 2  365

Tomczak-Jaegermann, N. See S. Mendelson Uspenskij, V. V. See E. Glasner Veličković, B. See B. König Wang, L. See X. Li